OpenStudy (vheah):

How to solve for t? For some reason im derping out and I'm forgetting how to deal with the cosine 0.090 = 0.180cos(0.989(t))

OpenStudy (518nad):

arccos or cos^-1 on your calculator

OpenStudy (518nad):

0.090 = 0.180cos(0.989(t)) 0.090/0.180=cos(0.989t) arccos(0.090/0.180) =0.989 t

OpenStudy (vheah):

Oh alright! i did that earlier but got the wrong answer. I must have plugged in the wrong numbers. thanks.

OpenStudy (518nad):

remember to use radians in this case

OpenStudy (vheah):

that might have bveen the problem because i think i left my calculator on degrees lol. thanks!

OpenStudy (518nad):

sure np :)